OpenStudy (anonymous):

a = ____, b = 24, c = 30

OpenStudy (anonymous):

a=18

OpenStudy (anonymous):

a^2 + 24^2 = 30^2 a^2 + 576 = 900 - 576 -576 a^2 = 324 a = 18